# Long-Context Autoregressive Video Modeling with Next-Frame Prediction

### v2

@@ Abstract
Long-context autoregressive modeling has significantly advanced language generation, but video generation still struggles to fully utilize extended temporal contexts. To investigate long-context video modeling, we introduce Frame AutoRegressive (FAR), a strong baseline for video autoregressive modeling. Just as language models learn causal dependencies between tokens (i.e., Token AR), FAR models temporal causal dependencies between continuous frames, achieving better convergence than Token AR and video diffusion transformers. Building on FAR, we observe that long-context video modeling faces challenges due to visual redundancy. Training on long videos is computationally expensive, as vision tokens grow much faster than language tokens. To tackle this issue, we propose balancing locality and long-range dependency through long short-term context modeling. A high-resolution short-term context window ensures fine-grained temporal consistency, while an unlimited long-term context window encodes long-range information using fewer tokens. With this approach, we can train on long video sequences with a manageable token context length, thereby significantly reducing training time and memory usage. Furthermore, we propose a multi-level KV cache designed to support the long short-term context modeling, which accelerating inference on long video sequences. We demonstrate that FAR achieves state-of-the-art performance in both short- and long-video generation, providing a simple yet effective baseline for video autoregressive modeling. The code is released at https://github.com/showlab/FAR.
@@ 1 Introduction
Advanced long-context autoregressive language models have demonstrated remarkable capabilities, enabling various applications that require test-time scaling, such as extended conversations [1], chain-of-thought reasoning [2, 3, 4], in-context learning [5], and retrieval-augmented generation [6]. However, video modeling has not achieved comparable progress. Recent video autoregressive modeling [7] directly adapts the paradigm of language models [8], where frames are factorized into discrete [9, 10] codes for next-token prediction (denoted as Token-AR). However, Token-AR still fails to achieve comparable quality to video diffusion transformers [11] due to the unidirectional modeling of visual tokens [12] and the irreparable information loss caused by vector quantization. On the other hand, video diffusion models [13, 14, 15, 16], which generate long videos using a progressive sliding window, struggle to effectively use earlier context.
@@
Figure 1: Evaluation on Long Video Prediction. FAR effective exploits long video contexts and achieves accurate prediction.
@@
In this paper, we introduce Frame AutoRegressive (FAR) model, specifically designed for video autoregressive modeling. FAR is trained using a frame-wise flow matching objective with autoregressive contexts. Unlike Token-AR, which learn causal dependencies between discrete tokens, FAR captures causal dependencies between continuous frames while still allowing full attention modeling within each frame. However, as a hybrid AR-Diffusion model [17, 18, 19], FAR also encounters a common issue observed in such models, namely, the discrepancy in observed contexts between training and inference. During training, later frames are exposed only to noised context frames due to diffusion objective, whereas inference relies on clean context frames. Recent methods [20, 21] mitigate this issue by appending a clean copy of the noised sequence during training, but this approach doubles the training cost.
To address the discrepancy of observed context, we propose training FAR with stochastic clean context. During training, we randomly replace a portion of noisy frames with clean frames and assign them a unique timestep embedding beyond the diffusion schedule to indicate representation extraction from clean context. During inference, this special embedding guides the model to effectively utilize clean context frames. We demonstrate FAR with stochastic clean context achieve same training efficiency to video diffusion transformers while achieving better convergence, served as a strong autoregressive video generation baseline.
Building on FAR, we investigate how to train on long video sequences to enable long-context video modeling. Unlike language modeling, long-context video modeling suffers from visual redundancy, as vision tokens grow significantly faster than language tokens when context increases. To address this challenge, we propose to balance the locality and long-range dependency with long short-term context modeling. Specifically, we maintain a high-resolution short-term context window to ensure fine-grained temporal consistency, while using a unlimited long-term context window with aggressive patchification to reduce redundant context tokens. This strategy enables efficient training on long video sequences with a manageable token context length. To speed up the proposed long short-term context modeling during inference, we introduce a multi-level KV cache to enhance inference efficiency.
@@
    Building on FAR, we introduce long short-term context modeling for efficient long-video training, as well as multi-level KV cache for faster inference.
@@
    FAR enables significantly lower training costs on long videos and achieves state-of-the-art performance in both short- and long-video generation.
@@ 2.2 Long-Context Language Modeling
A straightforward approach to improving long-context ability in language modeling is to directly fine-tune the model on longer sequences. Recent work [33, 34] has explored efficient long-sequence fine-tuning with position interpolation. However, training on long videos leads to prohibitive computational costs, as vision tokens scale much faster than language tokens with increasing context length. To address this issue, we introduce long short-term context modeling to reduce visual redundancy in long-video training.
@@ 2.3 Long-Context Video Modeling
Recent advancements in video generation models have enabled their use as interactive world simulators [35, 36, 37], which require the ability to exploit long-range context and memorize the observed environment. However, existing video diffusion transformers lack effective mechanism to utilize long-range context. Although early work [38] has explored long-context video prediction, it has been limited in visual quality and long-range consistency. In this work, we introduce FAR, a efficient framework for both short- and long-context autoregressive video modeling.
@@ 4 FAR
In this section, we first present the framework of FAR in Sec. 4.1. Then, we discuss the difficulties and solutions in training FAR in Sec. 4.2. In Sec. 4.3, we analyze the key design that enables FAR for long-context video modeling.
@@ 4.2 Short-Video Modeling
Our Solution: Stochastic Clean Context. To bridge the gap in observed context, we introduce stochastic clean context for training FAR. As illustrated in Fig. 2(a), we randomly replace a portion of the noised frames with their corresponding clean context and assign them a unique timestep embedding (e.g., -1) beyond the flow-matching timestep scheduler. These clean context frames are excluded from loss computation and are implicitly learned through later frames that use them as context. During inference, this unique timestep embedding guides the model to use clean context effectively.
Training FAR with stochastic clean context does not add extra computation and does not conflict with different timestep sampling strategies during training (e.g., logit-normal sampling [45]). It effectively resolves the training-inference discrepancy, as exemplified in Fig. 4(b).
@@ 4.3 Long-Context Video Modeling
Token Redundancy in Long Video. Visual data contains spatial redundancy, causing vision tokens to expand much faster than language tokens as context increases. For example, a video sequence of 128 frames requires more than 8K tokens, as illustrated in Fig. 7(a). Consequently, training on long videos becomes computationally prohibitive, as shown in Fig. 7(b, c).
Our Solution: Long Short-Term Context Modeling. To address the token redundancy in video, we introduce long short-term context modeling, which exploits the spatial and temporal locality in video data. As illustrated in Fig. 2(b), we maintain a high-resolution short-term context window to learn fine-grained temporal consistency and a low-resolution long-term context window, where we adopt aggressive patchification to reduce the number of context tokens. During training, given that the data has a maximum sequence length of $m$ frames, we fix the short-term context window to $n$ frames and randomly sample the long-term context frames from the range $[0,m-n]$. The attention mask with long short-term context modeling is shown in Fig. 3(b), where the long-term context uses fewer tokens. As demonstrated in Fig. 7(a), this strategy ensures that increasing the vision context length maintains a manageable token context length. With long short-term context modeling, we can reduce the cost and memory usage of the long-video training significantly, as shown in Fig. 7(b, c). To prevent interference between long-term and short-term contexts, we adopt separate projection layers for each context, inspired by MM-DiT [45]. This approach results in a slightly larger number of parameters, referred to as FAR-Long in Table. I.
TABLE II: Experimental Configurations of FAR. We follow the evaluation settings from Latte [44], MCVD [46], and TECO [38].
Hyperparameters Short-Video Generation Short-Video Prediction Long-Video Prediction
Cond. UCF-101 Uncond. UCF-101 BAIR UCF-101 Minecraft DMLab
Dataset Configuration
Resolution 256/128 256/128 64 64 128 64
Total Training Samples 13,320 13,320 43,264 9,624 194,051 39,375
Training Configuration
Batch Size 32 32 32 32 32 32
Latent Size 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47])
Training Sequence Length 16 16 32 16 300 300
LR $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
LR Schedule constant constant constant constant constant constant
Warmup Steps - - - - 10K 10K
Total Training Steps 400K 400K 200K 200K 1M 1M
Stochastic Clean Context 0.1 0.1 0.1 0.1 0.1 0.1
Short-Term Context Window 16 16 32 16 16 16
Long-Term Context Resolution - - - - 2$\times$2 2$\times$2
Evaluation Configuration
Samples 4$\times$2048 4$\times$2048 100$\times$256 100$\times$256 4$\times$256 4$\times$256
Guidance Scale 2.0 - - - 1.5 1.5
Reference Work Latte [44] Latte [44] MCVD [46] MCVD [46] TECO [38] TECO [38]
@@ 4.4 Inference-Time KV Cache
KV Cache for Short-Video Modeling. Due to the autoregressive nature of FAR, we can leverage KV-Cache to accelerate inference. As illustrated in Fig. 6, for each frame, we first use the flow-matching schedule to decode it into the clean latent frame. We then introduce an additional caching step to encode the clean latent frame into the KV cache. As discussed in Sec. 4.2, we use timestep $t=-1$ to denote the clean context frame in the caching step. These KV caches are subsequently used for autoregressive decoding of subsequent frames.
@@ 5.1 Implementation Details
We follow the DiT’s structure [39] to implement FAR. To compress video latents, we train a series of image DC-AE [47] on the corresponding dataset, resulting in 64 tokens per frame. All models are trained from scratch without image pretraining. We provide detailed training hyperparameters and evaluation setting in Table. II.
5.2 Quantitative Comparison
5.2.1 Video Generation
@@ 5.4 Ablation Study
Stochastic Clean Context. We have visualized the effectiveness of stochastic clean context in Fig. 4. Based on the quantitative evaluation of video prediction in Table. VI, FAR with stochastic clean context achieves significantly improved performance.
Multi-Level KV Cache. We evaluate the inference speed of FAR in Fig. 9. FAR without the KV cache incurs the highest inference cost, while the KV cache significantly reduces it. For long videos, FAR-Long with proposed multi-level KV cache achieves the lowest inference cost.
Long-Term Context Resolution. We investigate the impact of long-term context resolution on prediction accuracy and inference speed. As the context resolution increases, pixel-level metrics improve; however, the overall video quality remains similar. Nonetheless, inference speed significantly degrades at higher context resolutions due to the increased number of tokens involved in computation. Therefore, we select a 2$\times$2 resolution for the long-term context as a balance between computational efficiency and long-term context performance.
Short-Term Context Window Size. We evaluate the impact of the short-term context window size on performance. As shown in Fig. 11, video quality (FVD) quickly saturates as the short-term context window size increases, while pixel-level metrics continue to improve but also approach saturation at a window size of 16. Therefore, we set the short-term context window size to 16 by default.
TABLE VI: Ablation Study of Stochastic Clean Context on UCF-101. Stochastic clean context mitigates the training-inference discrepancy in observed context, leading to improved performance.
@@
FAR w/o. Stochastic Clean Context 0.540 16.42 0.211 399
FAR w/. Stochastic Clean Context 0.596 18.46 0.187 347
TABLE VII: Ablation Study on the Resolution of Long-Term Context. The speed is averaged over the generation of 300 frames.
@@ 6 Conclusion
In this paper, we introduce FAR for long-context video modeling. FAR effectively learns causal dependencies across continuous frames and demonstrates faster convergence compared to video diffusion transformers. To address visual redundancy in long-video training, we further propose a long short-term context modeling strategy, which encodes both remote and nearby visual contexts using asymmetric patchification. This approach enables a manageable token length as context frames increase, reducing both time and memory costs during long-video training. FAR achieves state-of-the-art performance on both short- and long-video modeling tasks, underscoring its potential as a new foundation model for video generation.

### v3

@@ Abstract
Long-context video modeling is essential for enabling generative models to function as world simulators, as they must maintain temporal coherence over extended time spans. However, most existing models are trained on short clips, limiting their ability to capture long-range dependencies, even with test-time extrapolation. While training directly on long videos is a natural solution, the rapid growth of vision tokens makes it computationally prohibitive. To support exploring efficient long-context video modeling, we first establish a strong autoregressive baseline called Frame AutoRegressive (FAR). FAR models temporal dependencies between continuous frames, converges faster than video diffusion transformers, and outperforms token-level autoregressive models. Based on this baseline, we observe context redundancy in video autoregression. Nearby frames are critical for maintaining temporal consistency, whereas distant frames primarily serve as context memory. To eliminate this redundancy, we propose the long short-term context modeling using asymmetric patchify kernels, which apply large kernels to distant frames to reduce redundant tokens, and standard kernels to local frames to preserve fine-grained detail. This significantly reduces the training cost of long videos. Our method achieves state-of-the-art results on both short and long video generation, providing an effective baseline for long-context autoregressive video modeling. The code is released at https://github.com/showlab/FAR.
@@ 1 Introduction
Long-context video modeling is essential for advancing video generative models toward real-world simulation [1]. However, current state-of-the-art video generative models (e.g., Wan [2], Cosmos [3]) fall short in this aspect. These models are typically trained on short video clips; for example, Wan and Cosmos are trained on approximately 5-second video segments. These methods effectively capture short-term temporal consistency, such as object or human motion. However, they fail to maintain long-term consistency, like memorizing the observed environments, which requires learning from long video observations.
Learning from long videos presents inherent challenges, primarily due to the prohibitive computational cost associated with processing the large number of vision tokens. As a result, many previous efforts have focused on test-time long video generation, employing training-free approaches to produce extended video sequences [4, 5]. However, while these methods can generate visually plausible long videos, they fail to effectively leverage long-range context. To truly capture long-range dependencies, it is essential to train or fine-tune models directly on long videos. However, this remains computationally expensive. Concurrent efforts, such as direct long-context tuning [6], face high computational costs, while test-time training [7] typically requires specialized architectural designs and incurs additional inference overhead. Therefore, an efficient framework for long-video training is urgently needed to enable effective long-context video modeling.
@@
Figure 1: Evaluation on Long Video Prediction. FAR better exploits long video contexts and achieves accurate prediction.
@@
To investigate this problem, we first establish a baseline for video autoregressive modeling, namely the Frame AutoRegressive (FAR) model. Unlike token-based autoregressive models (Token-AR), FAR operates in a continuous latent space, capturing causal dependencies between frames while still allowing full attention modeling within each frame. FAR is trained using a frame-wise flow matching objective with autoregressive context. As a hybrid AR-Diffusion model, FAR also encounters a gap between the observed context during training and inference. This is a common issue in similar models (e.g.,  [8], [9], [10]). To address this problem, we propose training FAR with stochastic clean context, which enables the model to leverage clean context signals during training and thus reduces bias at inference time. We demonstrate that FAR trained with stochastic clean context achieves better performance than video diffusion transformers and Token-AR, establishing itself as a strong autoregressive video generation baseline.
Building on FAR, we observe context redundancy in video autoregressive modeling. Specifically, the current frame relies more heavily on nearby frames to capture local motion consistency, while distant frames primarily function as context memory. To exploit this property, we introduce long short-term context modeling with asymmetric patchify kernels. In this approach, we apply a large patchify kernel to distant context frames in order to compress redundant tokens, while using the standard patchify kernel on nearby context frames to preserve fine-grained temporal consistency. This method significantly reduces training costs on long videos and leads to notable improvements in FAR’s long-context modeling capability, as demonstrated in action-conditioned long video prediction.
@@
    Building on FAR, we observe context redundancy in video autoregressive modeling and propose long short-term context modeling with asymmetric patchify kernels to substantially reduce long-video training costs.
@@
    FAR achieves state-of-the-art performance in both short- and long-video modeling.
@@ 2.2 Long-Context Language Modeling
Long-context language modeling typically follows two main approaches: test-time extrapolation and direct fine-tuning on long sequences. In the test-time extrapolation setting, many studies explore extrapolatable positional embeddings [28, 29, 30]. While these methods enable inference on longer sequences, they often underperform compared to models trained directly on long-text corpora. To reduce the computational cost, recent work [31, 32] has proposed efficient long-sequence fine-tuning strategies for large language models. However, training on long video sequences poses greater challenges, as vision tokens grow much faster than language tokens with increasing context length. To address this, we propose a long short-term context modeling approach using asymmetric patchify kernels, which effectively reduce context redundancy during long-video training.
@@ 2.3 Long-Context Video Modeling
Recent advancements in video generation models have enabled their use as interactive world simulators [33, 34, 35], which require the ability to exploit long-range context and memorize the observed environment. However, existing video diffusion transformers lack effective mechanism to utilize long-range context. Although early work [36, 37, 38] has explored long-video prediction, it has been limited in visual quality and long-range consistency. In this work, we introduce FAR, a efficient framework for long-context autoregressive video modeling.
@@ 4 FAR
We first present the FAR framework in Sec. 4.1, followed by training challenges and solutions in Sec. 4.2. Sec. 4.3 analyzes the design for long-context video modeling, and Sec. 4.4 introduces the KV cache for faster inference.
@@ 4.1 Framework Overview
TABLE II: Comparison of Routes for Long-Context Video Modeling. In test-time extrapolation, FAR is trained on short videos and evaluated on long videos using different extrapolation methods.
[…]
Test-Time Extrapolation
Sliding Window 0.365 12.3 0.415 161
Naive RoPE Ext. 0.372 12.2 0.397 396
RIFLEx [5] 0.372 12.2 0.398 391
Long-Video Training
FAR-B-Long 0.576 19.3 0.153 34
@@ 4.2 Short-Video Modeling
Our Solution: Stochastic Clean Context. To bridge the gap in observed context, we introduce stochastic clean context for training FAR. As illustrated in Fig. 2(a), we randomly replace a portion of the noised frames with their corresponding clean context and assign them a unique timestep embedding (e.g., -1) beyond the flow-matching timestep scheduler. These clean context frames are excluded from loss computation and are implicitly learned through later frames that use them as context. During inference, this unique timestep embedding guides the model to use clean context effectively. Training FAR with stochastic clean context does not add extra computation and does not conflict with different timestep sampling strategies during training (e.g., logit-normal sampling [45]). It effectively resolves the training-inference discrepancy, as exemplified in Fig. 4(b).
@@ 4.3 Long-Context Video Modeling
Test-Time Extrapolation vs. Long-Video Training. Two promising approaches to achieving long-context modeling in language modeling are test-time extrapolation [30, 29, 46] and long-sequence fine-tuning [31, 32]. In video modeling, most efforts [5, 15, 4] have focused on test-time extrapolation to generate long videos. However, we question whether test-time extrapolation can effectively solve long-context video modeling. To investigate this, we train FAR on action-conditioned short-video prediction and extend it to long-video prediction using various extrapolation techniques. As summarized in Table. II, test-time extrapolation results in significantly lower quality than the baseline sliding window approach, leading to poor predictions. Therefore, direct training on long videos may be necessary to achieve effective long-context video modeling.
Explosive Token Growth in Long Video Training. Visual data contains redundancy, causing vision tokens to expand much faster than language tokens as context increases. For example, a video sequence of 128 frames requires more than 8K tokens, as illustrated in Fig. 6(a). Consequently, training on long videos becomes computationally prohibitive, as shown in Fig. 6(b, c).
Our Solution: Long Short-Term Context Modeling. To address explosive token growth in long-video training, we leverage the concept of context redundancy. Specifically, in video autoregression, the current frame depends more heavily on nearby frames to capture local motion consistency, whereas earlier frames primarily serve as context memory. To eliminate context redundancy, we propose a long short-term context modeling with asymmetric patchify kernels. As shown in Fig. 2(b), we maintain a high-resolution short-term context window to capture fine-grained temporal consistency, and a low-resolution long-term context window, where we apply a large patchify kernel to compress context tokens. During training, given that the data has a maximum sequence length of $m$ frames, we fix the short-term context window to $n$ frames and randomly sample the long-term context frames from the range $[0,m-n]$. The attention mask with long short-term context modeling is shown in Fig. 3(b), where the long-term context uses fewer tokens. As demonstrated in Fig. 6(a), this strategy ensures that increasing the vision context length maintains a manageable token context length. With long short-term context modeling, we can reduce the cost and memory usage of the long-video training significantly, as shown in Fig. 6(b, c). To prevent interference between long-term and short-term contexts, we adopt separate projection layers for each context, inspired by MM-DiT [45]. This approach results in a slightly larger number of parameters, referred to as FAR-Long in Table. I.
@@ 4.4 Inference-Time KV Cache
KV Cache for Short-Video Modeling. Due to the autoregressive nature of FAR, we can leverage KV-Cache to accelerate inference. As illustrated in Fig. 7, for each frame, we first use the flow-matching schedule to decode it into the clean latent frame. We then introduce an additional caching step to encode the clean latent frame into the KV cache. As discussed in Sec. 4.2, we use timestep $t=-1$ to denote the clean context frame in the caching step. These KV caches are subsequently used for autoregressive decoding of future frames.
@@ 5.1 Implementation Details
We follow the DiT’s structure [39] to implement FAR. To compress video latents, we train a series of image DC-AE [47] on the corresponding dataset, resulting in 64 tokens per frame. All models are trained from scratch without image pretraining. We provide training hyperparameters and evaluation setting in the Table. VIII.
@@ 5.4 Ablation Study
How to decide patchify kernel for remote context? The key to selecting an appropriate patchify kernel lies in whether the compressed latent representation can reliably retain information from past observations. Since patchifying is a spatial-to-channel transformation, we hypothesize that it should satisfy the condition $c\times c\times d\leq D$, where $c$ is the patchify kernel size, $d$ is the latent dimension, and $D$ is the model’s channel dimension. For example, in our experiment, the latent dimension is 32 and the model dimension is 768. Using a patchify kernel of size $4\times 4$, we get $4\times 4\times 32=512<768$, which suggests that nearly all input information can be preserved during patchification. As demonstrated in Table. VII, the $[4,4]$ patchify kernel significantly reduces training cost, enabling feasible training on long videos, without sacrificing prediction accuracy compared to a larger $[8,8]$ kernel.
How to decide the local context length? We gradually increase the local context length during training and observe that the performance converges at certain short-term context length (i.e., 8 frames in Fig. 10). Further increasing the local context length significantly raises the training cost without improving performance. This experiment verifies the presence of context redundancy in video autoregressive modeling. Therefore, we select the saturate point as the optimal short-term context length.
TABLE VI: Ablation Study of Stochastic Clean Context (SCC) on UCF-101. Stochastic clean context mitigates the training-inference discrepancy in observed context, leading to improved performance.
@@
w/o. SCC 0.540 16.42 0.211 399
w/. SCC 0.596 18.46 0.187 347
Effect of Stochastic Clean Context. We have visualized the effectiveness of stochastic clean context in Fig. 4. Based on the quantitative evaluation of video prediction in Table. VI, FAR with stochastic clean context achieves significantly improved performance.
Effect of the KV Cache on Inference Speedup. As shown in  Fig. 11, the baseline FAR model, which samples without using long short-term context or KV cache requires approximately 1341 seconds to generate a 256-frame video. When the KV cache is introduced, the inference time is significantly reduced to 171 seconds. Finally, by incorporating both long short-term context and the corresponding multi-level KV cache, the sampling time further decreases to approximately 104 seconds. These results demonstrate that KV caching, especially when combined with long short-term context modeling, significantly improves sampling efficiency for long video generation.
TABLE VII: Ablation Study on the Patchify Kernel of Distant Context. Larger patchify kernels significantly reduce training cost.
@@ 6 Conclusion
In this paper, we systematically investigate long-context video modeling using the proposed Frame Autoregressive Model (FAR). First, We show that direct test-time extrapolation is insufficient for effective long-context video modeling, highlighting the necessity of efficient long-video training. We then identify context redundancy as a key bottleneck in video autoregression. To address this, we propose a long short-term context modeling strategy with asymmetric patchify kernels, a simple yet effective method to eliminate redundant context and significantly reduce the training cost for long videos. Extensive experiments validate the effectiveness of FAR in handling long-context video generation and highlight a promising direction for the evolution of next-generation video generative models—shifting the focus from short-term temporal consistency to long-term world modeling.